# Bi-Hamiltonian structures of WDVV-type

### v1

@@ 1.2 Bi-Hamiltonian structures for WDVV systems
Following [19], we introduce three new dependent variables $u^{1}=f_{xxx}$, $u^{2}=f_{txx}$, $u^{3}=f_{ttx}$. Then, the equation (5) can be rewritten in the form (18) as
@@ 1.3 Problem and results
Then, under the action of the full group of projective reciprocal transformations, it can be shown that all nontrivial cases reduce to one, the simplest WDVV equation. Unfortunately, the corresponding transformation is practically impossible to be found with current computational tools. This implies that our methods, which allow to find the bi-Hamiltonian structure, are still highly valuable.
@@ 1.4 Computational problems and their solutions
  • Finding a Hamiltonian formulation (2) through a first-order operator $P$ as in (3) amounts to solving a complicated system of nonlinear PDEs. Such a system is not easily solvable, in general, already when $n=3$.
@@ 3.1 Affine classification
$Q=\left(\begin{matrix}\gamma(c_{2}{-}c_{3})(u^{2})^{2}{+}(c_{1}\gamma u^{2}{-}c_{1})u^{1}{+}2c_{3}u^{2}&c_{2}(u^{1})^{2}{+}c_{1}u^{1}u^{2}{-}c_{3}(u^{2})^{2}{-}c_{3}\\ c_{2}(u^{1})^{2}{+}c_{1}u^{1}u^{2}{-}c_{3}(u^{2})^{2}{-}c_{3}&-c_{1}u^{1}+2c_{3}u^{2}+c_{3}\gamma\end{matrix}\right)$
@@ A Monge decomposition $(f^{(2)}_{ij})=\Psi\Phi\Psi^{\top}$ of the Monge metric $f^{(2)}$ is given by
$Q^{23}=-\mu^{2}(2BD+(u^{1}u^{2}-u^{3})(\alpha\delta-\beta\gamma))+4\mu(B^{2}+D^{2})-5BD-EF$
@@ 6 And beyond
$\begin{split}&u^{1}_{t}=\left(\frac{4(u^{1})^{3}u^{3}-4(u^{1})^{2}u^{6}+2u^{1}u^{2}u^{4}-(u^{4})^{2}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{2}_{t}=u^{1}_{x},\\ &u^{3}_{t}=\left(\frac{4(u^{1})^{2}(u^{3})^{2}-4u^{1}u^{3}u^{6}+u^{4}u^{5}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{4}_{t}=\left(\frac{2(u^{1})^{3}u^{5}-4(u^{1})^{2}u^{2}u^{6}+2(u^{1})^{2}u^{3}u^{4}+2u^{1}(u^{2})^{2}u^{4}-u^{2}(u^{4})^{2}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x}\\ &u^{5}_{t}=\left(\frac{2(u^{1})^{2}u^{3}u^{5}+2u^{1}(u^{3})^{2}u^{4}-2u^{1}u^{5}u^{6}+u^{2}u^{4}u^{5}-2u^{3}u^{4}u^{6}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x}\\ &u^{6}_{t}=\left(\frac{1}{2}u^{4}\right)_{x}\end{split}$ (18)

### v2

@@ 1.2 Bi-Hamiltonian structures for WDVV systems
Following [19], we introduce new dependent variables $u^{1}=f_{xxx}$, ${u^{2}=f_{txx}}$, $u^{3}=f_{ttx}$. Then, the equation (5) can be rewritten in the form (1) as
@@ 1.3 Problem and results
Then, under the action of the full group of projective reciprocal transformations, it can be shown that all nontrivial cases reduce to one, the simplest WDVV equation. Unfortunately, the corresponding transformation is practically impossible to find with current computational tools. This implies that our methods, which allow to find the bi-Hamiltonian structure, are still highly valuable.
@@ 1.4 Computational problems and their solutions
  • Finding a Hamiltonian formulation (2) through a first-order operator $P$ as in (3) amounts to solving a complicated system of nonlinear PDEs. Such a system is not easily solvable even when $n=3$.
@@ 3.1 Affine classification
$Q=\left(\begin{matrix}\gamma(c_{2}{-}c_{3})(u^{2})^{2}{+}(c_{1}\gamma u^{2}{-}c_{1})u^{1}{+}2c_{3}u^{2}&c_{2}(u^{1})^{2}{+}c_{1}u^{1}u^{2}{-}c_{3}(u^{2})^{2}{-}c_{3}\\ c_{2}(u^{1})^{2}{+}c_{1}u^{1}u^{2}{-}c_{3}(u^{2})^{2}{-}c_{3}&-c_{1}u^{1}+2c_{3}u^{2}+c_{3}\gamma\end{matrix}\right).$
@@ A Monge decomposition $(f^{(2)}_{ij})=\Psi\Phi\Psi^{\top}$ of the Monge metric $f^{(2)}$ is given by
$Q^{23}=-\mu^{2}(2BD+(u^{1}u^{2}-u^{3})(\alpha\delta-\beta\gamma))+4\mu(B^{2}+D^{2})-5BD-EF,$
@@ 6 And beyond
$\begin{split}&u^{1}_{t}=\left(\frac{4(u^{1})^{3}u^{3}-4(u^{1})^{2}u^{6}+2u^{1}u^{2}u^{4}-(u^{4})^{2}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{2}_{t}=u^{1}_{x},\\ &u^{3}_{t}=\left(\frac{4(u^{1})^{2}(u^{3})^{2}-4u^{1}u^{3}u^{6}+u^{4}u^{5}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{4}_{t}=\left(\frac{2(u^{1})^{3}u^{5}-4(u^{1})^{2}u^{2}u^{6}+2(u^{1})^{2}u^{3}u^{4}+2u^{1}(u^{2})^{2}u^{4}-u^{2}(u^{4})^{2}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{5}_{t}=\left(\frac{2(u^{1})^{2}u^{3}u^{5}+2u^{1}(u^{3})^{2}u^{4}-2u^{1}u^{5}u^{6}+u^{2}u^{4}u^{5}-2u^{3}u^{4}u^{6}}{2(2u^{1}u^{2}u^{3}-u^{1}u^{5}-u^{3}u^{4})}\right)_{x},\\ &u^{6}_{t}=\frac{1}{2}u^{4}_{x}.\end{split}$ (18)